OpenStudy (anonymous):

Please help, I have the answer for c but I don't know how to finish problem. Complete the square by finding the correct value for c. Then factor the resulting polynomial in to two binomials. x^2 + 18x + c c=81

OpenStudy (andrewhaze):

im not sure if its just me and you but i got the same

OpenStudy (anonymous):

yeah i know thats right, its the factoring part

OpenStudy (andrewhaze):

yup :)

OpenStudy (mathstudent55):

c = 81 is correct. \(a^2 + 2ab + b^2 = (a + b)^2\) \(a^2= x^2\), so \(a = x.\) \(b^2 = 81\), so \(b = 9\) \(x^2 + 18x + 81 = (x + 9)^2 = (x + 9)(x + 9)\)

OpenStudy (andrewhaze):

yes we already said that xD but thans for explanation better than min just plain out answer so for that you get a medal @mathstudent55

OpenStudy (andrewhaze):

me*

OpenStudy (anonymous):

thanks guys :)